OpenStudy (anonymous):

Solve the System of Equations 3x-4y=12 1/3x-4/9y=4/3

OpenStudy (anonymous):

y=3(x-4)/4 for the first one

OpenStudy (anonymous):

what is x and y =?

OpenStudy (owlcoffee):

So we have: 3x-4y=12 1/3x-4/9y=4/3 let's do it by substitution, meaning that we take one of the ecuations and solve it for a variable, I'll take the second one and solve it for y: \[\frac{ 1 }{ 3x-4 }-\frac{ 4 }{ 9y }=\frac{ 4 }{ 3 }\] Let's isolate y: \[\frac{ 1 }{ 3x-4 }=\frac{ 4 }{ 3 }+\frac{ 4 }{ 9y }\] \[\frac{ 1 }{ 3x-4 }=\frac{ 4(9y)+4(3) }{ 3(9y)}\]